# Multi-Field Adaptive Retrieval

### v1

@@ 3.3 Experimental Setup
mFAR affords a combination of lexical and dense scorers across experiments. Like our baselines, we use BM25 as our lexical scorer and Contriever as our dense scorer. Because of potential differences across datasets, we initially consider four configurations that take advantage of mFAR’s ability to accommodate multiple fields or scorers: mFAR Dense uses all fields and the dense scorer, mFAR Lexical uses all fields and the lexical scorer, mFAR All uses all fields and both scorers, and mFAR 2 uses both scorers but the single-field (Sec. 3.1) document representation. Based on the initial results, we additionally consider mFAR 1+n, which consists of a single-field lexical scorer and multi-field dense scorer. This results in models that use $|\mathcal{F}|$, $|\mathcal{F}|$, $2|\mathcal{F}|$, $2$, and $|\mathcal{F}|+1$ scorers respectively. For each model and each dataset, we run a grid search over learning rates and whether to normalize and select the best model based on the development set.
Because Contriever is based on a 512-token window, we prioritize maximizing this window size for each field, which ultimately reduces the batch size we can select for each dataset. We use effective batch sizes of 96 for Amazon and Prime, and 192 for MAG, and train on 8xA100 GPUs. We use patience-based early stopping based on the validation loss. More details on the exact hyperparameters for each run are in Appendix B.
Table 2: Comparing our method (mFAR) against baselines and state-of-the-art methods on the STaRK test sets. ada-002 and multi-ada-002 are based on vector similarity; +{Claude3, GPT4} further adds an LLM reranking step on top of ada-002. AvaTaR is an agent-based iterative framework. Contriever-FT is a finetuned Contriever model, which is also the encoder finetuned in mFAR. mFAR is superior against prior methods and datasets, and earns a substantial margin on average across the benchmark. * In Wu et al. (2024b), these are only run on a random 10% subset.
@@
Claude3 reranker* 0.455 0.538 0.559 0.365 0.484 0.442 0.178 0.356 0.263 0.333 0.459 0.421
GPT4 reranker* 0.448 0.554 0.557 0.409 0.486 0.490 0.183 0.341 0.266 0.347 0.460 0.465
AvaTaR (agent) 0.499 0.606 0.587 0.444 0.506 0.512 0.184 0.393 0.267 0.376 0.502 0.455
@@ 4 Results
We report the results from our mFAR models in Table 2, compared against against prior methods and baselines. Our best models—both make use of both scorers—perform significantly better than prior work and baselines: mFAR 2 on Amazon and MAG, and mFAR All on Prime and STaRK average. This includes surpassing re-ranking based methods and the strongest agentic method, AvaTaR. mFAR All performs particularly well on Prime (+20% for H@$1$). Comparatively, all models based on ada-002 have extended context windows of 2K tokens, but mFAR, with an encoder that has a much smaller context window size (512), still performs significantly better. Furthermore, our gains cannot be only attributed to finetuning or full reliance on lexical scorers since the mFAR models perform better against the already competitive BM25 and finetuned Contriever baselines.
@@
A side-by-side comparison of the single-field models against their multi-field counterparts shows mixed results. If we only consider dense scorers, mFAR Dense produces the better results than Contriever-FT across all datasets. To our knowledge, this is the first positive evidence in favor of multi-field methods in dense retrieval. For mFAR Lexical, in both Amazon and MAG, the BM25 baseline performs especially well, and we do not see consistent improvements. This specific phenomenon has been previously noted by Robertson et al. (2004), who describe a modification that should be made to the BM25 algorithm itself to properly aggregate multi-field information. Specifically in BM25, the scores are length normalized. For some terms, like institution, repetition does not imply a stronger match, and so treating the institution field separately (and predicting high weights for it) could lead to high scores for negative documents. A multi-field sparse representation, then, may not always be the best solution, depending on the dataset.
@@
Across both multi-field (mFAR All vs. mFAR Dense or mFAR Lexical) and single-field models (mFAR 2 vs. BM25 or Contriever-FT), and across almost every dataset, there is an increase in performance when using both scorers over a single scorer type, validating our earlier hypothesis. This reinforces findings from prior work (Gao et al., 2021; Kuzi et al., 2020) that hybrid methods work well. The one exception (Prime, single-field) may be challenging for single-field models, possibly due to the relatively higher number of fields in the dataset and the semantics of the fields, as we investigate more in Section 5.3. However, in the multi-field setting for Prime, we again see hybrid perform best. This provides evidence for our original motivation: that hybrid models are suitable for and positively benefits certain structured, multi-field documents.
Combining the above findings: that multi-field dense retrieval is superior, single-field BM25 can be more effective, and that hybrid works best, we experiment with mFAR 1+n, a single-field lexical, multi-field dense model. This obtains the highest average scores, including the highest score on Amazon and MAG.
@@ 5 Analysis
5.1 Is our query-conditioned adaptation necessary?
@@
Table 3: The test scores of mFAR All without query conditioning (QC) and the % lost without it.
@@ 5.2 Which fields and scorers are important?
The interpretable design of our mFAR framework enables us to easily control the used fields and scorers after a model has been trained. Specifically, we can mask (zero) out any subset of the weights $w_{f}^{m}$ used to compute $s(q,d)$ (Equation 4). For example, setting $w_{f}^{\text{lexical}}=0$ for each $f$ would force the model to only use the dense scores for each field. We can interpret a drop in performance as a direct result of excluding certain fields or scorers, and thus we can measure their contribution (or lack thereof). In this deep-dive analysis, we study mFAR All and re-evaluate the model’s performance on each dataset after masking out entire scoring methods (lexical or dense), specific fields (title, abstract, etc), and even specific field and scoring method (e.g. title with dense scorer).
@@
By performing similar analysis at a fine-grained field-level, we can identify which parts of the document are asked about or useful. For each field $f_{i}$, we can set $w_{f_{i}}^{\text{lexical}}=0$, $w_{f_{i}}^{\text{dense}}=0$, or both. We collect a few interesting fields from each dataset in Table 5, with all fields in Appendix D.
@@
To simplify the model, one may suggest removing any $s_{f}^{m}(q,d)$ where setting $w_{f}^{m}=0$ results in no drop. However, we cannot do this without hurting the model. In other words, low deltas do not imply low importance. For some fields (e.g. Amazon’s qa, MAG’s title, or Prime’s phenotype absent), when the lexical or dense scorers are zeroed out individually, the scores are largely unaffected. However, completely removing the field by zeroing both types of scorers results in a noticeable drop. In many cases, we observe that masking out entire fields yields a larger drop than masking out either one individually. This type of behavior could be a result of mFAR redundantly obtaining the same similarity information using different scorers. On the contrary, there is also information overlap across fields, and so in some cases, it is possible in some cases to remove entire fields, especially in Prime (e.g. enzyme) and Amazon, without substantial drops.
Table 5: For each dataset, the absolute change (delta) of masking out certain fields and scorers from mFAR All for H@$1$ and R@$20$. For each field, we zero out either the lexical scorer, the dense scorer, or both. The raw scores on all metrics for all fields in each dataset are in Appendix D.
@@ 5.3 Qualitative Analysis
In Figure 3 (top), the query is looking for either a gene or protein that is not expressed. With mFAR All, the retriever matches a longer text more accurately than the single-field hybrid retriever doesours Hybridand mFAR 2 correctly match female gonad. However, mFAR selects the field that refers to the absence of an expression, which the model learns. In the mFAR 2, because the lexical scorer cannot distinguish between present and absent and the document embedding also failed to do so, it incorrectly ranks a negative document higher.
@@ 6 Related Work
Multi-field and Hybrid Retrieval.
Multi-field retrieval has previously been explored in the context of sparse methods (Robertson et al., 2004; Zaragoza et al., 2004) which can involve extensions of the lexical-based BM25 (Robertson et al., 1994), learned sparse representations (Zamani et al., 2018), or Bayesian approaches (Piwowarski & Gallinari, 2003). Recently, Lin et al. (2023) approached a similar task with a primary focus on query decomposition and rewriting. They also propose a weighted combination of “expert” retrievers, though their weights are hand-picked and each (dense) retriever is trained independently. In mFAR, we finetune the entire system end-to-end with a single shared encoder and our weights are learned, which allows us to scale to dozens of fields.
The combination of both types of lexical and dense scorers has previously been found to be complementary, leading to performance gains (Gao et al., 2021; Kuzi et al., 2020). Notably, Kuzi et al. (2020) points out that long documents are challenging for lexical-based methods and suggests document chunking as a possible remedy in future work. We implicitly segment the document by taking advantage of the multi-field structure inherently present in documents, and unlike those past works, our work is the first to demonstrate the strength of hybrid-based methods in a multi-field setting.
While we focus on documents with multi-field structure, prior work has focused on documents with other types of structure. For example, Chen et al. (2020b) design a model for table search, while Husain et al. (2019) propose the task for code search. Reddy et al. (2022) introduce an Amazon shopping dataset with short search queries and proposes an embeddings-based method for document ranking, relevance classification, and product substitutions. We already use an Amazon dataset within STaRK, include several embeddings-based baselines, and are more interested in queries that relate to multiple fields, which is not the focus of their task. The dataset from Lin et al. (2023) is multimodal; our work is unimodal: we focus first on natural language and next the question of lexical and dense scorers.
Some prior works have incorporated document structure as part of pretraining. Su et al. (2023) generates pseudo-queries based on Wikipedia formatting, while Li et al. (2023) introduces an alignment objective to increase similarity between the structured part of the dataset and its corresponding natural-language description. Both of these pretraining-focused methods assume some redundancy in the underlying data, which STaRK, and multi-field documents in general, do not guarantee, and we focus on introducing a post-training method that does not modify the underlying model pretraining scheme.
@@ 7 Conclusion
We present mFAR, a novel framework for retrieval over multi-field data by using multiple scorers, each independently scoring the query against a part (field) of a structured document. These scorers can be lexical-based or dense-based, and each field can be scored by both types. We introduce an interpretable and controllable query-conditioned predictor of weights used to adaptively sum over these scores. On three large-scale datasets, we find that mFAR can achieve significant performance gains over existing methods due to multi-field advantages and the inclusion of a hybrid combination of scorers, leading to state-of-the-art performance. Through our analysis, we find that the best models benefit from both access to both scorers and the ability to weight each field conditioned on the query, further verifying our method.
Our primary goal is to study the challenging and emerging problem of retrieval for multi-field structured data and to introduce a flexible framework to approach it. Having laid the groundwork, future work can include more specialized individual scorers, scale up to more scorers in other modalities like vision or audio, and add other algorithmic improvements to the weighted integration of scores across scorers. Then, mFAR would be a step towards retrieval of any type of content, which can further aid applications for general search or agent-based RAG.
@@ A.1 Preprocessing
Technically, STaRK is a dataset of queries over knowledge graphs. The baselines, Wu et al. (2024b), create a linearized document for each node, which omits some edge and multi-hop information that is available in the knowledge graph. AvaTaR (Wu et al., 2024a) operates directly on the knowledge graph. As we want to operate over structured documents, we need a preprocessing step either on the linearized documents or by processing the graph.
@@ A.2 Fields
We include the list of fields that we used in this work in Table 6. Not every single field available in the STaRK knowledge graph (Wu et al., 2024b) is used because some are not used in the baseline and we try to match the baselines as closely as possible. We make some cosmetic changes for space and clarity in the examples in the main body of this paper, including uppercasing field names and replacing underscore with spaces. We also shorten “author___affiliated_with___institution”, “paper___cites___paper” to “Papers Cited”, “paper___has_topic___field_of_study” to “Area of Study” and expand “type” to “Entity Type.”
@@ Appendix B Implementation Details
During training, we sample 1 negative example per document. Along with in-batch negatives, this results in $12b-$ 1 negative samples for a batch size of $b$. This negative document is sampled using Pyserini Lucene [^4] https://github.com/castorini/pyserini: 100 nearest documents are retrieved, of which the postive documents are removed. The top negative document is then sampled among that remaining set. We apply early stopping on validation loss with a patience of 5. We set $\tau$ = 0.05 and train with DDP on 8x NVIDIA A100s. Contriever is a 110M parameter model, and the additional parameters added through $G$ is negligible ($768|\mathcal{F}|$), scaling linearly in the number of fields.
We use separate learning rates (LRs) for finetuning the encoder and for the other parameters. Specifically, we searched over learning rates [5e-6, 1e-5, 5e-5, 1e-4] for the encoder and [1e-3, 5e-3, 1e-2, 5e-2, 1e-1] for the parameters in $G(q,f,m)$ which consist of $\mathbf{a}_{f}^{m}$ and $\gamma_{f}^{m}$ and $\beta_{f}^{m}$ from batch normalization. The main grid search was conducted over the bolded values, although we found 5e-3 to be effective for $G(q,f,m)$ for Amazon. We otherwise follow the default settings for both the optimizer (AdamW, dropout, etc. from sentence-transformers 2.2.2) and batch normalization (PyTorch 2.4.0). As mentioned, whether to apply batch normalization at all was also a hyperparameter searched over: we found it useful in the hybrid setting. Our implementation uses Pytorch Lightning [^5] https://lightning.ai/ and sentence-transformers (Reimers & Gurevych, 2019). We use a fast, python-based implementation of BM25 as our lexical scorer (Lù, 2024). [^6] https://github.com/xhluca/bm25s The best hyperparameters for each of our models in this work are listed in Table 7. In the case where there is only a single field (last two sections), the adaptive query conditioning is not needed.
@@
Model Dataset Referenced in Encoder LR $G()$ LR Batch norm?
mFAR All Amazon Table 2, 3, most tables/figures 1e-5 5e-3 no
@@
mFAR 1+n Amazon Table 2 1e-5 1e-2 no
@@
Appendix C Full Results on STaRK

### v2

@@ 3.3 Experimental Setup
mFAR affords a combination of lexical and dense scorers across experiments. Similarly to our baselines, we use BM25 as our lexical scorer and the dot product of Contriever embeddings as our dense scorer. We use a shared embedding model for both the query and document and for creating $\mathbf{q}$ when computing the adaptation function $G$. Because of potential differences across datasets, we initially consider four configurations that take advantage of mFAR’s ability to accommodate multiple fields or scorers: mFAR Dense uses all fields and the dense scorer, mFAR Lexical uses all fields and the lexical scorer, mFAR All uses all fields and both scorers, and mFAR 2 uses both scorers but the single-field (Sec. 3.1) document representation. Based on our final results and analysis, we additionally create and evaluate mFAR All+2, which consists of both a single-document and multi-field representation for both lexical and dense scoring methods. This results in five mFAR models that use $|\mathcal{F}|$, $|\mathcal{F}|$, $2|\mathcal{F}|$, $2$, and $22|\mathcal{F}|+$ 2 scorers respectively. For each dataset (and across models), we run a grid search over learning rates and whether to normalize and select the best model based on the development set.
Because Contriever has a 512-token context window, we prioritize maximizing this window size for each field, which ultimately reduces the batch size we can select for each dataset, resulting in 96 for Amazon and Prime, and 192 for MAG. More details on the exact hyperparameters for each run are in Appendix B.
[…]
@@
Claude3 0.455 0.538 0.559 0.365 0.484 0.442 0.178 0.356 0.263 0.333 0.459 0.421
GPT4 0.448 0.554 0.557 0.409 0.486 0.490 0.183 0.341 0.266 0.347 0.460 0.465
AvaTaR agent 0.499 0.606 0.587 0.444 0.506 0.512 0.184 0.393 0.267 0.376 0.502 0.455
@@ 4 Results
We report the results from our mFAR models in Table 1, compared against against prior methods and baselines. Our best models—both make use of both scorers—perform significantly better than prior work and baselines: mFAR 2 on Amazon, and mFAR All and mFAR All+2 on the other datasets. This includes surpassing re-ranking based methods and the strongest agentic method, AvaTaR. mFAR All performs particularly well on Prime (+20% for H@$1$). Comparatively, all models based on ada-002 have extended context windows of 2K tokens, but mFAR, using an encoder that has a much smaller context window size (512), still performs significantly better. Furthermore, our gains cannot be only attributed to finetuning or full reliance on lexical scorers since the mFAR models perform better against the already competitive BM25 and finetuned Contriever baselines.
@@
A side-by-side comparison of the single-field models against their multi-field counterparts shows mixed results. If we only consider dense scorers, mFAR Dense produces the better results than Contriever-FT across all datasets. To our knowledge, this is the first positive evidence in favor of multi-field methods in dense retrieval. For mFAR Lexical, in both Amazon and MAG, the BM25 baseline performs especially well, and we do not see consistent improvements. This specific phenomenon has been previously noted by Robertson et al. (2004), who describe BM25F, a variant of BM25 that aggregates multi-field information in a more principled manner. Specifically in BM25, the scores are length normalized. For some fields, like institution, repetition does not imply a stronger match, and so treating the institution field separately (and predicting high weights for it) could lead to high scores for negative documents. A multi-field sparse representation, then, may not always be the best solution, depending on the dataset. We further try using BM25F instead but find lackluster performance likely due to undertuned weights (Appendix D). Given the number of fields in STaRK datasets, tuning field weights is a challenging and open problem that appears less tractable for BM25F than for mFAR. Finally, we note that combining multi-field with single-field can lead to further gains, as demonstrated by mFAR All+2 (and Appendix C.2).
@@
Across both multi-field (mFAR All vs. mFAR Dense or mFAR Lexical) and single-field models (mFAR 2 vs. BM25 or Contriever-FT), and across almost every dataset, there is an increase in performance when using both scorers over a single scorer type, validating our earlier hypothesis. This reinforces findings from prior work (Gao et al., 2021; Kuzi et al., 2020) that hybrid methods work well. The one exception (Prime, single-field) may be challenging for single-field models, possibly due to the relatively higher number of fields in the dataset and the semantics of the fields, as we investigate more in Section 5.3. However, in the multi-field setting for Prime, we again see hybrid perform best. This provides evidence for our original motivation: that hybrid models are suitable for and positively benefits certain semi-structured, multi-field documents.
@@ 5 Analysis
Our analyses, along with the quantitative results, lead us to experiment with a combination of single-field and multi-field document representations in Appendix C.2, like mFAR All+2. We find that these combinations offer additional gains over mFAR All, and that even just a combination of single-field lexical and multi-field dense improves over only mFAR All or mFAR 2.
5.1 Is query-conditioned adaptation necessary?
@@
Table 2: The test scores of mFAR All without query conditioning (QC) and the % relative change without it.
@@ 5.2 Which fields and scorers are important?
The interpretable design of our mFAR framework enables us to easily control the used fields and scorers after a model has been trained. Specifically, we can mask (zero) out any subset of the weights $w_{f}^{m}$ used to compute $s(q,d)$ (Equation 4). For example, setting $w_{f}^{\text{lexical}}=0$ for each $f$ would force the model to only use the dense scores for each field. We can interpret a drop in performance as a direct result of excluding certain fields or scorers, and thus we can measure their contribution (or lack thereof). In this deep-dive analysis, we re-evaluate mFAR All’s performance on each dataset after masking out entire scoring methods (lexical or dense), specific fields (title, abstract, etc), and even specific field and scoring method (e.g. title with dense scorer).
@@
By performing similar analysis at a fine-grained field-level, we can identify which parts of the document are asked about or useful. For each field $f_{i}$, we can set $w_{f_{i}}^{\text{lexical}}=0$, $w_{f_{i}}^{\text{dense}}=0$, or both. We collect a few interesting fields from each dataset into Table 4, with all fields in Appendix E.
@@
To simplify the model, one may suggest removing any $s_{f}^{m}(q,d)$ where setting $w_{f}^{m}=0$ results in no drop. However, we cannot do this without hurting the model. In other words, low deltas do not signify low importance. For some fields (e.g. Amazon’s qa, MAG’s title, or Prime’s phenotype absent), when the lexical or dense scorers are zeroed out individually, the scores are largely unaffected. However, completely removing the field by zeroing both types of scorers results in a noticeable drop. In many cases, we observe that masking out entire fields yields a larger drop than masking out either one individually. This type of behavior could be a result of mFAR redundantly obtaining the same similarity information using different scorers. On the contrary, there is also information overlap across fields, and so in some cases, it is possible in some cases to remove entire fields, especially in Prime (e.g. enzyme) and Amazon, without substantial drops.
Table 4: For each dataset, the absolute change (delta) of masking out certain fields and scorers from mFAR All for H@$1$ and R@$20$. For each field, we zero out either the lexical scorer, the dense scorer, or both. The raw scores on all metrics for all fields in each dataset are in Appendix E.
@@ 5.3 Qualitative Analysis
In Figure 3 (top), the query is looking for either a gene or protein that is not expressed. With mFAR All, the retriever matches a longer text more accurately than mFAR 2 does. Both mFAR All and mFAR 2 correctly match female gonad. However, mFAR All selects the field that refers to the absence of an expression, which is learned by the model. In mFAR 2, because the lexical scorer cannot distinguish between present and absent, mFAR 2 incorrectly ranks the negative document higher.
@@ 6 Related Work
Structured and Semi-structured Retrieval
Forms of structured and semi-structured retrieval have been explored in a variety of tasks and domains. In particular, we focus on multi-field retrieval, a form of semi-structured retrieval (Zaragoza et al., 2004) for which prior sparse approaches include the aforementioned BM25F (Robertson et al., 2004), learned sparse representations (Zamani et al., 2018), and Bayesian approaches (Piwowarski & Gallinari, 2003). Lin et al. (2023) approach a similar task using dense retrievers and with a primary focus on query decomposition to support a weighted combination of “expert” retrievers. In contrast to their system in which weights are hand-picked and each retriever is independently trained, our mFAR model is learned end-to-end with a single shared encoder and learned weights, which aids scalability.
Table retrieval (Zhang & Balog, 2021; Bhagavatula et al., 2015; Pasupat & Liang, 2015; Herzig et al., 2021; Shraga et al., 2020; Chen et al., 2020b) is a structured retrieval task which adopts similar methods as multi-field retrieval (see Appendix F for an evaluation of mFAR for this task). Table retrieval does not necessarily require table-specific model design, as linearized forms of the table can be adequate for competitive performance (Wang et al., 2022) and many table retrieval datasets have been seen during encoder pretraining (e.g. DPR (Karpukhin et al., 2020) has been trained on Wikipedia). Beyond tabular data, other structured retrieval tasks include code search (Husain et al., 2020) and knowledge graph datasets like shopping (Reddy et al., 2022). The latter is similar to the Amazon subset of STaRK. Separately, a books QA dataset used by Lin et al. (2023) is multimodal, which is not our focus.
Besides decomposition, which targets parts of the structure, like fields, with specialized parameters, prior work has also investigated modifying the training process through generating pseudo-queries based on Wikipedia formatting (Su et al., 2024) and incorporating auxiliary alignment objectives between the document and a natural language description (Li et al., 2023). These methods generally assume that there exist semantically repetitive information (e.g. table and description of table). We do not make this assumption and focus on post-training methods that can use off-the-shelf encoders.
Hybrid Methods
The combination of both types of lexical and dense scorers has previously been found to be complementary, leading to performance gains (Gao et al., 2021; Kuzi et al., 2020; Lee et al., 2023). Notably, Kuzi et al. (2020) points out that long documents are challenging for lexical-based methods and suggests document chunking as a possible remedy in future work. We implicitly segment the document by taking advantage of the multi-field structure inherently present in documents, and unlike those past works, our work is the first to demonstrate the strength of hybrid-based methods in a multi-field setting. Alternative hybrid retrieval setups combine both dense and lexical features with a dense encoder, trained end-to-end (Lin & Lin, 2023; Shen et al., 2023), whereas we explicitly use existing lexical scorers.
@@ 7 Conclusion
We present mFAR, a novel framework for retrieval over multi-field data by using multiple scorers, each independently scoring the query against a part (field) of a semi-structured document. These scorers can be lexical-based or dense-based, and each field can be scored by both types. We introduce an interpretable and controllable query-conditioned predictor of weights used to adaptively sum over these scores. On three large-scale datasets, we find that mFAR can achieve significant performance gains over existing methods due to multi-field advantages and the inclusion of a hybrid combination of scorers, leading to state-of-the-art performance. Through our analysis, we find that the best models benefit from both access to both scorers and the ability to weight each field conditioned on the query, further verifying our method.
Our primary goal is to study the challenging and emerging problem of retrieval for multi-field semi-structured data and to introduce a flexible framework to approach it. Having laid the groundwork, future work can include more specialized individual scorers, scale up to more scorers in other modalities like vision or audio, and add other algorithmic improvements to the weighted integration of scores across scorers. Then, mFAR would be a step towards retrieval of any type of content, which can further aid applications for general search or agent-based RAG.
8 Acknowledgements
The authors thank the reviewers for the engaging exchange and their helpful feedback. The authors also thank Nick Craswell, Jason Eisner, Jenny Liang, Shirley Wu, Jay DeYoung, Somin Wadhwa, and Byron Wallace, for the insightful discussions to improve the paper. ML conducted this work as a research intern in the MSAI group at Microsoft. ML is also supported by an NSF Graduate Research Fellowship.
@@ Appendix A Dataset
Dataset Domain Num. Documents Num. Fields Train Dev. Test.
Amazon products, product reviews 950K 8 6K 1.5K 1.5K
MAG science papers, authors 700K 5 8K 2.6K 2.6K
Prime biomedical entities 130K 22 6.1K 2.2K 2.8K
@@ A.1 Preprocessing
Technically, STaRK is a dataset of queries over knowledge graphs. The full dataset details are in Table 5. The baselines (Wu et al., 2024b) create a linearized document for each node, which omits some edge and multi-hop information that is available in the knowledge graph. AvaTaR (Wu et al., 2024a) operates directly on the knowledge graph. As we want to operate over semi-structured documents, we need a preprocessing step either on the linearized documents or by processing the graph.
@@ A.2 Fields
We include the list of fields that we used in this work in Table 6. Not every single field available in the STaRK knowledge graph (Wu et al., 2024b) is used because some are not used in the baseline and so we try to match the baselines as closely as possible. We make some cosmetic changes for space and clarity in the examples in the main body of this paper, including uppercasing field names and replacing underscore with spaces. We also shorten “author___affiliated_with___institution”, “paper___cites___paper” to “Papers Cited”, “paper___has_topic___field_of_study” to “Area of Study” and expand “type” to “Entity Type.”
@@ Appendix B Implementation Details
During training, we sample $k=1$ negative example per query. Along with in-batch negatives, this results in $12b-$ 1 negative samples for a batch size of $b$. This negative document is sampled using Pyserini Lucene [^4] https://github.com/castorini/pyserini: 100 nearest documents are retrieved, of which the postive documents are removed. The top negative document is then sampled among that remaining set. We apply early stopping on validation loss with a patience of 5. We set $\tau$ = 0.05 and train with DDP on 8x NVIDIA A100s. Contriever is a 110M parameter model, and the additional parameters added through $G$ is negligible ($768|\mathcal{F}|$), scaling linearly in the number of fields.
We use separate learning rates (LRs) for finetuning the encoder and for the other parameters. Specifically, we searched over learning rates [5e-6, 1e-5, 5e-5, 1e-4] for the encoder and [1e-3, 5e-3, 1e-2, 5e-2, 1e-1] for the parameters in $G(q,f,m)$ which consist of $\mathbf{a}_{f}^{m}$ and $\gamma_{f}^{m}$ and $\beta_{f}^{m}$ from batch normalization. The main grid search was conducted over the bolded values, although we found 5e-3 to be effective for $G(q,f,m)$ for Amazon. We otherwise follow the default settings for both the optimizer (AdamW, dropout, etc.) and batch normalization (PyTorch 2.4.0). As mentioned, whether to apply batch normalization at all was also a hyperparameter searched over: we found it useful in the hybrid setting.
Our implementation uses Pytorch Lightning [^5] https://lightning.ai/ and sentence-transformers 2.2.2 (Reimers & Gurevych, 2019). We use a fast, python-based implementation of BM25 as our lexical scorer (Lù, 2024). [^6] https://github.com/xhluca/bm25s The best hyperparameters for each of our models in this work are listed in Table 7. In the case where there is only a single field (last two sections), the adaptive query conditioning is not needed.
@@
Model Dataset Mainly referenced in Encoder LR $G()$ LR Batch norm?
mFAR All Amazon Table 1, most tables/figures 1e-5 5e-3 no
@@
mFAR Dense&1 Amazon Table 10 (appendix only) 1e-5 1e-2 no
@@
Appendix C Detailed Results on STaRK